# Survey of Graph Neural Network for Internet of Things and NextG Networks

### v1

@@ V-E Digital Twin Networks
In [222], the authors propose TwinNet, a GNN-based DT model that can take into account the complex relationship between different service level agreements metrics such as queuing policy, network topology, routing configuration, and input traffic. It is shown that GNN-enabled TwinNet can generalize to its input parameters, topologies, routing and queuing configuration, and can accurately predict the end-to-end path delays in 106 unseen real-world topologies from Internet Topology Zoo [226]. Flow-aware Digital Twin (FlowDT) [223], is a GNN-based framework to model computer networks, where flow refers to the set of packets that share common characteristics such as the same protocols (UDP/TCP), IP addresses or port numbers of source and destination node. FlowDT supports the creation and destruction of flows and also provide accurate estimates of per-flow metrics such as delay and jitter.
V-F Unmanned Aerial Networks
Unmanned Aerial Networks (UANs) refer to networks formed by interconnected Unmanned Aerial Vehicles (UAVs), also known as drones, operating cooperatively to achieve various objectives. Due to the portable nature of the UAVs, they can be dynamically deployed when required and play a crucial role in establishing temporary connectivity in the event of disaster or infrastructure failure and enable high data rate network connectivity. Figures 11 and 12 show the Typical components of UANs and applications of GNN for UAN tasks, respectively. ML techniques such as RNNs and RL are commonly used for UAN problems such as spectrum access and flight control for microwave, millimeterwave, and Terahertz frequency bands [227, 228, 229, 230, 231, 232], UAV swarm collaborative beamforming [233, 234, 235, 236]. In addition to traditional ML techniques GNNs are also commonly used for UANs problems such as Sensing [237] and networking applications such as trajectory planning [238, 239], resilient UAV swarm communication [240], Flying Relay Location and Routing Optimization [241].
In [237], the authors propose a GCN and GAT-based LSTM model to accurately predict the real-time and future Air Quality Index (AQI) using aerial-ground sensing framework. This model uses a lightweight Dense-MobileNet model [242] to achieve energy-efficient end-to-end learning from features extracted from images captured by the UAV. This framework allows for federation across various institutions to collaboratively learn a well-trained global model to accurately and securely monitor the AQI and also expanding the scope of UAV swarm monitoring. In [238], the authors present a GCN-based framework to solve the cooperative trajectory design problem formulated based on modeling the relationship between the ground terminal and UAV-base stations (UAV-BS). This framework enables UAV-BS to efficiently manage the time-varying local observation such as current location, average throughput, and scheduling and facilitate cooperation between UAV-BSs through information exchange. Due to the lack of global information and a globally optimal solution, a Q-learning-based multi-agent RL algorithm is used by the UAV-BS to learn a distributed trajectory policy which is modeled as a Markov game. Here, the message passing by the GCN framework ensures cooperative behavior among UAV-BS and significantly enhances the network’s performance. In [239], the authors propose a multi-dimensional trajectory prediction model for UAV swarms by integrating GNN with a dynamic graph neural network (DynGN). The DynGN uses an encode-decoder structure with a GCN to process the dynamic adjacency matrix, extract spatial features, and gated recurrent unit to model the temporal dynamics and capture the evolution of UAV interactions over time. This model processes both the dynamically changing network configuration and the trajectory data of the UAV swarm simultaneously.
In [240], the authors study the self-healing problem of UAV swarm network to re-establish network connectivity in the event of unpredictable destruction. These destructions may affect one or more UAVs simultaneously and can either be one-off events or recurring events. To address the one-off events, a GCN-based framework is designed to find the recovery topology of the UAV swarm network in a timely manner. This framework is then extended by incorporating a monitoring mechanism to detect the UAV destruction and a self-healing trajectory planning algorithm based on GCN to address the recurring events of destruction. Finally, a meta-learning scheme is designed for the GCN to reduce the execution time complexity and ensure faster convergence of GCN. In [241], the authors propose a GNN-based framework for the optimization of UAV locations and relay paths in UAV-relayed 6G IoT networks. A two-stage GNN model is developed to decouple the problem of optimal relay path selection and UAV position optimization and is solved by a separate GNN model. Here, both stages are trained using reinforcement and unsupervised learning with no training data or prior knowledge of optimal solutions. First, the locations of the UAVs are optimized and second, the best relay path is determined based on the optimized location of the UAVs from the first step. Results show that the proposed model performs better than brute search with lower time complexity, scalable to very large networks with dynamically changing environments.
@@ VI GNN for Tactical Systems
Tactical systems play a crucial role in modern warfare by enhancing the effectiveness, survivability, of military forces operating in dynamic and challenging environments. It is important to maintain a safe communication system and to identify and localize any imminent threat. To this end, in this section, we discuss the application of GNN for Tactical Communication and Sensing Systems (Section VI-A), Target Recognition (Section VI-B) and Localization (Section VI-C). Figure 13 shows applications of GNN for tactical system tasks.
@@ VI-A Tactical Communication and Sensing Systems
In [243], the authors propose context-surrounding graph neural networks with numbers (CS-GNN-N) for knowledge-based reasoning for military decision support. One of the major components is GNN-based learning which utilizes two smoothness metrics, namely feature smoothness and label smoothness to measure the quantity and quality of neighborhood information of nodes. In [244], the authors present a GNN-based mission planning approach for unmanned ground vehicles (UGV) to generate optimal mission planning results using recursive graph grammar to describe the task assignment process of multiple UGVs as an irregular graph structure and by coupling both task assignment as well as path planning. The performance of the proposed algorithm was verified both using simulation and real-world UGV robot. In [245], the authors propose GraphNearbyAlert (GNA) framework to effectively broadcast cooperative awareness messages to a group of vehicles. The results show that it is sufficient to transmit very little information about the topology of the graph to extract information from the shared data and can be used for decision making in cooperative transportation systems while preserving privacy. In [246] the authors present a GCN framework integrated with deep Q network based RL-based to achieve an efficient and reliable multi-agent decision making system for safe and efficient operation of connected autonomous vehicles in intelligent transportation systems. The GCN generates the features of the network based on the topology and is used by the RL to generate optimal decisions. In [247], the authors present an improved single-agent GNN-based algorithm for decision making in autonomous vehicles, where the GNN is first used to generate features for a multi-agent system and then transferred to a single-agent system. The advantage of this approach is that the newly trained single-agent model can then be used to train scenarios with multiple agents. This design significantly reduces the time complexity of training a multi-agent model and improves the convergence speed of the model.
@@
In [248], the authors introduce a novel GNN based behavior prediction model to train and develop synthetic characters for roles in the Rapid Integration and Development Environment (RIDE). RIDE is a unity-based military training simulation environment that facilitates rapid development and prototyping of simulated environments in direct service of the Army and other Department of Defense entities. Through simulation experiments, the authors show that a GNN model trained using pseudo-human behavior can be used as a behavior predictor and help with better cooperation between human agents. In [249], the authors propose a GNN-based DRL framework to autonomously accomplish real-time resource dispatching and allow human intervention to accomplish complex tactics. In this design, the GNN is used to generate an integrated scheduling model of detection, tracking and interception and transformed into a sequential decision problem. The GNN model is integrated with the proximal policy optimization (PPO) algorithm to learn the air defense environment. This framework enables human-intelligence collaborative dynamic scheduling for emergency response scenarios. In [250], the authors propose a multi-modal adaptive STGN framework to simultaneously explore spatio-temporal dependencies in the airspace network with the objective to effectively learn the diverse spatial relationship and adaptively adjust the impact of different spatial models. The results show that the framework can harness differing spatial models and achieve high generalization performance across different temporal patterns compared to state-of-the-art methods. In [251], the authors present a GCN-based multi-agent DRL routing algorithm for multi-sink tactical mobile sensor networks to overcome link-layer jamming attacks. The proposed algorithm captures the hop count to the nearest sink (i.e., ), the one-hop delay, the next hop’s packet loss, and the energy cost of packet forwarding. The results show that the proposed scheme outperforms compared to other state-of-the-art algorithms.
@@ VI-B Target Recognition
In addition to its use for decision-making, GNN has widely been used for automatic target recognition. Synthetic Aperture Radar (SAR) images are commonly utilized in military applications for automatic target recognition (ATR). SAR is a remote sensing technology that uses radar to create high-resolution images of the Earth’s surface. SAR works by emitting radio waves at a target and then measuring the reflection of those waves. Unlike optical imaging, SAR can operate day or night and in various weather conditions because it doesn’t rely on visible light. Due to this ability, SAR images are widely used in the military for ATR [252, 253, 254]. In addition to SAR images there are also other techniques used for target recognition such as Hyperspectral images [255, 256] and Infrared thermal imaging [257]. In [252], authors use GNN to identify ground-based objects such as battle tanks, personnel carriers, and missile launchers, among others. By identifying the type of objects, it would allow the military to determine if the objects are friendly or adversarial. However, the classification is susceptible to the noise level in the SAR images which may degrade the performance of the GNN classifier. The effect of an adversarial attack on GNN-based SAR image classifier has been extensively in [253]. Furthermore, extraction of features from the original SAR image requires a large amount of data and is time-consuming. To overcome this issue, in [254], the authors use SAR image target pixel gray-scale decline by a graph representation. Here, the gray-scale SAR image is divided into subintervals and a node is assigned to represent each pixel with declined order of pixel gray-scale in the subinterval. The raw SAR image is then transformed to construct a graph structure which is used by the GCN to extract the features of the graph structured data and then classify the target.
@@ VI-C Localization
Localization refers to the process of determining the location or position of an object or entity within a given environment. It is a fundamental task in various fields, including robotics, navigation, mobile computing, and wireless communication. Accurate localization is crucial for applications such as GPS-Based Navigation, Augmented Reality (AR), and autonomous vehicles. GNNs are used for localization in certain contexts due to their ability to capture spatial relationships and dependencies among different elements in a graph-based representation of the environment. GNNs can integrate information from different types of sensors or data sources within the same graph structure. This is beneficial in scenarios where multiple sensors (e.g., GPS, Wi-Fi, LIDAR, cameras) are available for localization. Localization data can be noisy and uncertain due to factors like signal interference, multi-path effects, and environmental changes. GNNs can learn to handle such noise and uncertainty by aggregating information from neighboring nodes and making robust predictions.
@@
GNNs are widely used for indoor localization [259] [260], large-scale network localization [261] and false data injection attack localization [262]. For example, in [259] graphs are generated using the RSSI from Wi-Fi access points to localize the device into certain predefined areas using GCN. Similarly, in [260], Zhang et al. proposed a novel WiFi domain adversarial graph convolutional network model for indoor localization using RSSIs between waypoints and WiFi access points. The difference in the two works is that in the latter unlabeled data and multiple data domains are considered and a semi-supervised domain adversarial training scheme was used to efficiently use the unlabeled data as well as different data distributions across data domains. While GNNs are widely used for indoor localization, they also find application in large-scale network localization such as [261] where a network with 500 nodes was considered in comparison to 25 nodes/devices in [259]. In [262], the authors propose a GNN-based approach to identify the presence of a false data injection attack as well as localize the source of the attack.

### v2

@@ V-F Unmanned Aerial Networks
V-G Lessons Learned
In this section, we present the key takeaways and lessons learned from the applications of GNN for Networking discussed in Sections V-A - V-F.
  1. 1.
    Despite their advantages, GNNs face scalability challenges as network size increases. The computational complexity of GNNs grows with the number of nodes and edges, making it difficult to apply them directly to large-scale networks like the Internet or hyper-scale data centers. Large networks require extensive message passing between nodes, leading to increased memory and processing requirements. Techniques such as sampling [301, 302], and distributed GNN architectures [303, 304] can be explored to mitigate these scalability constraints.
  2. 2.
    Networking operates across multiple layers of the OSI model, and performance optimization often requires a holistic, cross-layer approach. GNNs have the potential to model interactions between different layers, such as how physical-layer signal interference affects network-layer routing decisions. However, most research in GNN-based networking has focused on single-layer optimization, such as routing or traffic prediction. The exploration of multi-layer GNN models could lead to breakthroughs in end-to-end network performance optimization, including adaptive modulation, dynamic bandwidth allocation, and interference-aware routing.
  3. 3.
    UAV networks such as drone swarms and airborne communication relays, introduce unique challenges due to their rapidly changing topologies. Traditional static network models struggle to adapt to such dynamic environments. GNNs, with their ability to process graph structures dynamically, can model UAV network topologies in real-time, enabling efficient routing, link prediction, and connectivity management. They can be particularly useful for optimizing communication in ad hoc UAV networks where maintaining connectivity is a key challenge.
  4. 4.
    In conclusion, GNNs represent a transformative approach to network management and optimization, offering powerful capabilities in understanding and predicting complex network behaviors. Their ability to model network topologies, adapt to dynamic environments, and optimize multiple layers simultaneously makes them highly attractive for applications such as traffic engineering, anomaly detection, and resource allocation. However, significant challenges remain, particularly in scaling GNNs to large networks, ensuring real-time processing, and designing energy-efficient architectures. Addressing these challenges through novel algorithmic and hardware advancements will be key to unlocking the full potential of GNNs in next-generation networking. As research in this domain progresses, GNNs are poised to play a crucial role in building intelligent, resilient, and self-optimizing networks of the future.
@@ VI GNN for Tactical Systems
Modern warfare and defense operations rely on tactical systems to ensure effective coordination, secure communication, and resilient decision-making in highly dynamic and adversarial environments. Unlike conventional communication systems, tactical networks must adapt in real time, detect and respond to threats, and ensure operational integrity despite interference or adversarial attacks. These requirements extend beyond the capabilities of traditional NextG networks, necessitating intelligent, resilient, and adaptive solutions. Tactical systems rely on interdependent components such as distributed sensors, mobile units, and adversarial entities which is hard to be captured and modelled using traditional NNs but GNNs can be used effectively to model these complex interactions. To this end, in this section, we discuss the application of GNN for Tactical Communication and Sensing Systems (Section VI-A), Target Recognition (Section VI-B), and Localization (Section VI-C). Figure 16 shows applications of GNN for tactical system tasks. Tactical systems are an extension of NextG networks where it is not enough to be aware of the spectrum and improve the QoS, but also be able to take countermeasures to detect, and localize any nodes or adversary causing interference or affecting the integrity of the NextG networks. Table VIII provides a high-level comparison of surveyed paper using GNNs for tactical applications. We also provide the dataset or the simulator used for performance evaluation.
TABLE VIII: Comparison of surveyed papers using GNN for Tactical Systems Applications.
Section Paper Application ML Algorithm Used Dataset/Simulator Used
Section VI-A Tactical Communication and Sensing Systems Nie et al. [305] Military decision support Context-surrounding GNN with numbers DBP15K [306], own Milassistant dataset
Xu et al. [307] Mission planning GNN Own python-based simulator
Sant’Ana da Silva et al. [308] Decision making in intelligent transportation systems GNN FB15k [309]
Liu et al. [310] Decision-making in interactive traffic scenarios GNN + RL Own python-based simulator
Yang et al. [311] Decision-making in autonomous driving GNN + RL SUMO simulator [312]
Liu et al. [313] Behavior Prediction GNN + RL RIDE [314], OpenAI Gym [249]
Tang et al. [315] Real-time resource dispatching GNN + RL Own simulator
Li et al. [316] Airspace complexity prediction Spatio-temporal graph neural network Chinese domestic airspace traffic data
Okine et al. [317] Tactical wireless sensor networks GNN + RL Own python-based simulator
Section VI-B Target Recognition Wijeratne et al. [318] Target recognition GraphSAGE + CNN MSTAR [319]
Ye et al. [320] SAR image classifier GNN MSTAR [319]
Zhu et al. [321] Target classification Graph Convolutional Network MSTAR [319]
Xu et al. [322] Hyperspectral target detection GNN + CNN HYDICE dataset, AVRIS San Diego Dataset [323], Airport Beach Urban Dataset [324]
Chen et al. [325] Target detection Graph Convolution Networks + Graph Transformer AVRIS San Diego Dataset [323], Urban Dataset [324], MUUFL Gulfport Dataset [326]
Yang et al. [327] Aerial target pose estimation Motion flow-based Graph Convolutional Network Microsoft Visual Studio air combat simulation platform
Section VI-C Localization Lezama et al. [328] Wi-Fi fingerprinting based indoor localization Graph Convolutional Network MNAV [329] ; UJIIndooeLoc [330]
Zhang et al. [331] WiFi indoor localization Domain adversarial graph convolutional network Microsoft Indoor Location Competition 2.0 data set [332]
Yan et al. [125] Large-scale network localization Graph Convolutional Network Own simulator [333]
Boyaci et al. [334] Detection and localization of stealth false data injection attacks Auto-Regressive Moving Average Graph Filters-based GNN NYISO [335]
@@ VI-A Tactical Communication and Sensing Systems
In [305], the authors propose context-surrounding graph neural networks with numbers (CS-GNN-N) for knowledge-based reasoning for military decision support. One of the major components is GNN-based learning which utilizes two smoothness metrics, namely feature smoothness and label smoothness to measure the quantity and quality of neighborhood information of nodes. In [307], the authors present a GNN-based mission planning approach for unmanned ground vehicles (UGV) to generate optimal mission planning results using recursive graph grammar to describe the task assignment process of multiple UGVs as an irregular graph structure and by coupling both task assignment as well as path planning. The performance of the proposed algorithm was verified both using simulation and real-world UGV robots. In [308], the authors propose GraphNearbyAlert (GNA) framework to effectively broadcast cooperative awareness messages to a group of vehicles. The results show that it is sufficient to transmit very little information about the topology of the graph to extract information from the shared data and can be used for decision-making in cooperative transportation systems while preserving privacy. In [310] the authors present a GCN framework integrated with deep Q network-based RL-based to achieve an efficient and reliable multi-agent decision-making system for safe and efficient operation of connected autonomous vehicles in intelligent transportation systems. The GCN generates the features of the network based on the topology and is used by the RL to generate optimal decisions. In [311], the authors present an improved single-agent GNN-based algorithm for decision-making in autonomous vehicles, where the GNN is first used to generate features for a multi-agent system and then transferred to a single-agent system. The advantage of this approach is that the newly trained single-agent model can then be used to train scenarios with multiple agents. This design significantly reduces the time complexity of training a multi-agent model and improves the convergence speed of the model.
@@
In [313], the authors introduce a novel GNN-based behavior prediction model to train and develop synthetic characters for roles in the Rapid Integration and Development Environment (RIDE). RIDE is a unity-based military training simulation environment that facilitates rapid development and prototyping of simulated environments in direct service of the Army and other Department of Defense entities. Through simulation experiments, the authors show that a GNN model trained using pseudo-human behavior can be used as a behavior predictor and help with better cooperation between human agents. In [315], the authors propose a GNN-based DRL framework to autonomously accomplish real-time resource dispatching and allow human intervention to accomplish complex tactics. In this design, the GNN is used to generate an integrated scheduling model of detection, tracking, and interception and transformed into a sequential decision problem. The GNN model is integrated with the proximal policy optimization (PPO) algorithm to learn the air defense environment. This framework enables human-intelligence collaborative dynamic scheduling for emergency response scenarios. In [316], the authors propose a multi-modal adaptive STGN framework to simultaneously explore spatio-temporal dependencies in the airspace network with the objective to effectively learn the diverse spatial relationship and adaptively adjust the impact of different spatial models. The results show that the framework can harness differing spatial models and achieve high generalization performance across different temporal patterns compared to state-of-the-art methods. In [317], the authors present a GNN-based multi-agent DRL routing algorithm for multi-sink tactical mobile sensor networks to overcome link-layer jamming attacks. The proposed algorithm captures the hop count to the nearest sink (i.e., ), the one-hop delay, the next hop’s packet loss, and the energy cost of packet forwarding. The results show that the proposed scheme outperforms compared to other state-of-the-art algorithms.
@@ VI-B Target Recognition
While Tactical systems discussed in the previous section help in making critical decisions and mission planning, it is also important to manage and keep track of deployed resources. These resources can either be personnel or deployed ground and aerial resources. In this section, we will discuss several applications where GNNs are commonly used for target recognition. In addition to its use for decision-making, GNN has widely been used for automatic target recognition. Synthetic Aperture Radar (SAR) images are commonly utilized in military applications for automatic target recognition (ATR). SAR is a remote sensing technology that uses radar to create high-resolution images of the Earth’s surface. SAR works by emitting radio waves at a target and then measuring the reflection of those waves. Unlike optical imaging, SAR can operate day or night and in various weather conditions because it doesn’t rely on visible light. Due to this ability, SAR images are widely used in the military for ATR [318, 320, 321]. In addition to SAR images there are also other techniques used for target recognition such as Hyperspectral images [322, 325] and Infrared thermal imaging [327]. In [318], authors use GNN to identify ground-based objects such as battle tanks, personnel carriers, and missile launchers, among others. By identifying the type of objects, it would allow the military to determine if the objects are friendly or adversarial. However, the classification is susceptible to the noise level in the SAR images which may degrade the performance of the GNN classifier. The effect of an adversarial attack on GNN-based SAR image classifier has been studied extensively in [320]. Furthermore, extraction of features from the original SAR image requires a large amount of data and is time-consuming. To overcome this issue, in [321], the authors use SAR image target pixel gray-scale decline by a graph representation. Here, the gray-scale SAR image is divided into sub-intervals and a node is assigned to represent each pixel with declined order of pixel gray-scale in the sub-interval. The raw SAR image is then transformed to construct a graph structure which is used by the GCN to extract the features of the graph structured data and then classify the target.
@@ VI-C Localization
In the previous section, we discussed the application of GNN for target recognition. The obvious next step is to accurately pinpoint the location of the deployed resources as well as keep track of the movement through a process called localization. Localization refers to the process of determining the location or position of an object or entity within a given environment. It is a fundamental task in various fields, including robotics, navigation, mobile computing, and wireless communication. Accurate localization is crucial for applications such as GPS-Based Navigation, Augmented Reality (AR), and autonomous vehicles. GNNs are used for localization in certain contexts due to their ability to capture spatial relationships and dependencies among different elements in a graph-based representation of the environment. GNNs can integrate information from different types of sensors or data sources within the same graph structure. This is beneficial in scenarios where multiple sensors (e.g., GPS, Wi-Fi, LIDAR, cameras) are available for localization. Localization data can be noisy and uncertain due to factors like signal interference, multi-path effects, and environmental changes. GNNs can learn to handle such noise and uncertainty by aggregating information from neighboring nodes and making robust predictions.
@@
GNNs are widely used for indoor localization [328] [331], large-scale network localization [125] and false data injection attack localization [334]. For example, in [328] graphs are generated using the RSSI from Wi-Fi access points to localize the device into certain predefined areas using GCN. Similarly, in [331], Zhang et al. proposed a novel WiFi domain adversarial graph convolutional network model for indoor localization using RSSIs between way-points and WiFi access points. The difference in the two works is that in the latter unlabeled data and multiple data domains are considered. It also uses a semi-supervised domain adversarial training to efficiently use the unlabeled data as well as different data distributions across data domains. While GNNs are widely used for indoor localization, they also find application in large-scale network localization such as [125] where a network with 500 nodes was considered in comparison to 25 nodes/devices in [328]. In [334], the authors propose a GNN-based approach to identify the presence of a false data injection attack as well as localize the source of the attack.
VI-D Lessons Learned
In this section, we present the key takeaways and lessons learned from the applications of GNN for Tactical Systems discussed in Sections VI-A - VI-C.
  1. 1.
    Tactical environments, such as battlefields and military operations, involve highly dynamic and complex interactions among various entities, including soldiers, vehicles, and drones. These interactions are inherently interconnected, as communication and coordination between units play a crucial role in mission success. GNNs provide enhanced situational awareness, enabling military units to make more informed decisions. By accurately capturing the evolving state of the tactical environment, GNNs assist in mission planning, unit coordination, and threat assessment, ultimately improving operational effectiveness.
  2. 2.
    In battlefield scenarios, target recognition is often complicated by ambiguous or cluttered surroundings, where multiple objects can be closely positioned or occluded. Traditional object detection methods may struggle to differentiate between targets and non-targets in such scenarios. GNNs improve target recognition by leveraging spatial and relational information between detected objects. By considering contextual relationships such as the expected proximity of friendly and enemy assets, GNNs can refine detection accuracy and reduce false positives or negatives. This capability is particularly beneficial in surveillance and reconnaissance tasks, where precision and rapid identification are critical.
  3. 3.
    Effective localization and tracking of assets and targets are essential in tactical operations. These tasks often involve multiple sensors, including GPS, radar, and communication nodes, which must collaboratively determine real-time positions. GNNs improve localization accuracy even in GPS-denied environments [337, 338, 339]. This ensures that commanders maintain accurate knowledge of friendly and enemy unit positions, thereby optimizing resource deployment and mission execution.
  4. 4.
    Modern military operations require seamless interoperability across multiple domains, including land, sea, air, space, and cyber. Traditional communication systems often struggle to integrate information from these diverse domains due to differences in data formats, protocols, and communication architectures. GNNs facilitate cross-domain interoperability by modeling interactions between heterogeneous data sources within a unified graph framework. This capability allows for coordinated decision-making across different operational theaters, ensuring that forces can act cohesively even in highly distributed environments.
  5. 5.
    Training and simulation play a vital role in preparing military personnel for real-world operations. GNNs can be used to model various tactical scenarios, allowing commanders and troops to simulate battlefield conditions and assess different strategies. By analyzing past mission data and generating realistic engagement models, GNN-based simulations provide an advanced training environment where personnel can refine their decision-making skills under diverse operational conditions. This leads to better preparedness and adaptability in actual combat situations.
  6. 6.
    In conclusion, GNNs presents a great opportunity for enhancing tactical communication systems by providing superior modeling and real-time decision-making abilities. By integrating GNNs thoughtfully within tactical systems, military forces can achieve enhanced coordination, strategic advantage, and operational success in dynamic and high-stakes environments.